# On the First Passage Times of Branching Random Walks in $\mathbb{R}^{d}$

### v1

@@ Abstract
We study the first passage times of discrete-time branching random walks in ${\mathbb{R}}^{d}$ where $d\geqslant 1$. Here, the genealogy of the particles follows a supercritical Galton-Watson process. We provide asymptotics of the first passage times to a ball of radius one with a distance $x$ from the origin, conditioned upon survival. We provide explicitly the linear dominating term and the logarithmic correction term as a function of $x$. The asymptotics are precise up to an order of $o_{\mathbb{P}}(\log x)$ for general jump distributions and up to $O_{\mathbb{P}}(\log\log x)$ for spherically symmetric jumps. A crucial ingredient of both results is the tightness of first passage times. We also discuss an extension of the first passage time analysis to a modified branching random walk model that has been proven to successfully capture shortest path statistics in polymer networks.
@@ 1 Introduction
We consider a discrete-time branching random walk (BRW) on $\mathbb{R}^{d}$, where particles in $\mathbb{R}^{d}$ perform independent random walks until they randomly branch into more particles or die. Spatial branching processes, including BRW and its continuous-time sibling branching Brownian motion (BBM), have a long history and a wide range of applications in ecology, population biology, and modeling epidemics ([25, 28, 34, 35, 36, 49]). The study of the extremal behavior of spatial branching processes has gained increasing attention over recent years. The mathematical techniques applied in those studies are related to developments in numerous other fields, e.g., discrete Gaussian free field [14, 24], random unitary matrices [3], Riemann zeta function [4], random multiplicative functions [31], spin glasses [12], among many others that exhibit a log-correlated structure. We refer to [46, 52] for lecture notes on BRW and related topics.
@@
The study of the extrema for one-dimensional BBM traces back to [17], which established a precise asymptotic for the maximum as a function of time. For general branching random walks, [1] provided a precise asymptotic in terms of the jump distribution of the process. Finer behavior near the frontier is also well understood since the works of [2, 13, 38, 39]. Nevertheless, less is known in dimensions greater than one. To name a few examples, there are studies of the maximum norm for BBM ([7, 32, 33, 40, 47]) and very recently of BRW ([8]), as well as the range of spatial branching processes ([37, 53]). These works in higher dimensions typically assumed that the increment distribution is spherically symmetric. A notable exception is [48], which provided estimates on the number of particles in a linearly shifted ball in time for a continuous-time version of the BRW.
Our primary focus is on the first passage times (FPT) for branching random walks in $\mathbb{R}^{d}$. By definition, the FPT is the first time for particles to reach a prescribed region. For BBM, [54] established a precise asymptotic formula (up to $O_{\mathbb{P}}(1)$) for the FPT; see (62) below. Our interest in estimating the FPT of spatial branching processes stems from their relevance to polymer physics. The mechanical behavior of polymer networks appears to be controlled by the shortest paths between distant nodes in the polymer network, as shown by recent works [50, 51] using coarse-grained molecular dynamics (CGMD) simulations. The shortest path length (SP) serves as an indicator of how stretched the average load-bearing chains are in the polymer network. Theories based on SP go beyond traditional polymer models focused on elastic behaviors and aim to explain the inelastic behaviors exhibited by such materials.
The SP between (fixed) distant nodes is equivalent to the FPT of the BRW beginning from the origin (source node) and terminating at a desired point in space (destination node). Moreover, the occurrence of the cross-links along the polymer backbone mimics a Poisson process. As a result, a cross-link between two chains is equivalent to a branching event observed in the BRW genealogy tree. The agreement of the mean FPT from the BRW theory with the mean SP from the CGMD simulation results has been shown in our recent work [54]. Furthermore, the numerical implementation of the BRW model shows an excellent agreement between its FPT distribution and the SP distribution from the CGMD simulation. Our earlier work initially focused on constant length jumps with the BRW model where there existed a correlation between the components of a jump step in the $d$ dimensions by virtue of being sampled from the unit sphere $\mathbb{S}^{d-1}=\{\mathbf{x}\in\mathbb{R}^{d}\mid\left\lVert\mathbf{x}\right\rVert=1\}$. This was later shown to be represented equally well by the Gaussian BRW and BBM in which the components of each jump step are independent in the $d$ dimensions. However, the analysis was restricted to spherically symmetric jumps to accurately model the as-prepared polymer network prior to deformation.
@@
In this paper, we provide FPT asymptotics for BRW that may not have a spherically symmetric jump distribution (while the case of spherically symmetric jumps is of special interest for our applications). The main results are explained in Section 1.2 and their proofs are presented in Section 2. Applications to polymer physics will be detailed in Section 3.1.
@@ 1.1 Notation and assumptions
Let us start with a formal definition of the branching random walk (BRW). We consider a discrete-time BRW model with offspring distribution $\{p_{i}\}_{i\geqslant 0}$. This means that at each time step, each particle independently reproduces $i$ particles at the same location with probability $p_{i}$. The case of $i=0$ corresponds to the particle being terminated. Let $\rho=\sum_{i}i\,p_{i}$ be the mean of the offspring distribution. We will always assume the supercritical case of $\rho>1$. The $d$-dimensional random walk increment is given by ${\bm{\xi}}$, and the first particle starts from the origin $\mathbf{0}\in\mathbb{R}^{d}$. Typically, in this paper, a bold symbol refers to a vector.
Our goal is to study the asymptotic of the FPTs of the BRW. For $x\in\mathbb{R}$, we let $B_{x}$ denote a ball of radius one [^1] Replacing the radius one by any fixed positive constant will not change the analysis; the same proof goes through. centered at $(x,0,\dots,0)$ in $\mathbb{R}^{d}$. We let $V_{n}$ denote the collection (i.e. set) of particle locations at time step $n$, and $\{\bm{\eta}_{v,n}(k)\}_{0\leqslant k\leqslant n}$ denotes the $d$-dimensional random walk that leads to $v\in V_{n}$. We define the FPT $\tau_{x}$ of the BRW to $B_{x}$, that is,
$\tau_{x}=\min\{n\geqslant 0:\exists\,v\in V_{n},\leavevmode\nobreak\ \bm{\eta}_{v,n}(n)\in B_{x}\}.$
In what follows, we will separately analyze the cases depending on whether the law of ${\bm{\xi}}$ is spherically symmetric or not. [^2] A probability distribution on $\mathbb{R}^{d}$ is spherically symmetric if it is invariant under any orthonormal transformation. Equivalent definitions can be found in Theorem 2.5 of [26]. In both cases, we impose the following assumptions on the BRW:
@@
For a stochastic process $\{T_{x}\}_{x\geqslant 0}$ and a positive deterministic sequence $\{t_{x}\}_{x\geqslant 0}$, we write $T_{x}=O_{\mathbb{P}}(t_{x})$ if the collection $\{T_{x}/t_{x}\}_{x\geqslant 0}$ is tight, and $T_{x}=o_{\mathbb{P}}(t_{x})$ if $T_{x}/t_{x}\to 0$ in probability.
@@ 1.1.1 Spherically symmetric jumps
Denote the first coordinate of ${\bm{\xi}}$ by $\xi$, which is a real-valued random variable. We introduce the large deviation rate function
$I(x):=\sup_{\lambda>0}\Big{(}\lambda x-\log\phi_{\xi}(\lambda)\Big{)},$ (1)
where $\phi_{\xi}(\lambda):=\mathbb{E}[e^{\lambda\xi}]$ is the moment generating function for $\xi$. In this case, we make the following assumptions on ${\bm{\xi}}$:
@@
    the law of ${\bm{\xi}}$ is spherically symmetric in $\mathbb{R}^{d}$, and its radial component $R$ satisfies $\mathbb{P}(R=0)<1$;
@@
    $\log\rho\in(\mathrm{ran}I)^{\circ}$, where $(\mathrm{ran}I)^{\circ}$ is the interior of the range of $I$. In other words, there exists $c_{1}>0$ such that $I(c_{1})=\log\rho$. It can be shown that $c_{1}\in(\mathrm{ran}(\log\phi_{\xi})^{\prime})^{\circ}$.
[…]
@@ 1.1.2 Non-spherically symmetric jumps
Let
@@
    there exist $m\in\mathbb{N}$ and $r>1$ such that ${\bm{\xi}}_{1}+\dots+{\bm{\xi}}_{m}$ has a density in $L^{r}(\mathbb{R}^{d})$, where $\{{\bm{\xi}}_{i}\}_{1\leqslant i\leqslant m}$ are i.i.d. copies of ${\bm{\xi}}$;
@@
    $\log\rho\in(\mathrm{ran}I(\cdot,\mathbf{0}))^{\circ}$, where $I(\cdot,\mathbf{0})$ refers to the function $I$ with the last $d-1$ variables fixed at zero; let $\widehat{c}_{1}>0$ satisfy $I(\widehat{c}_{1},\mathbf{0})=\log\rho$, then $(\widehat{c}_{1},\mathbf{0})\in(\mathrm{ran}\nabla\log\phi_{\bm{\xi}})^{\circ}$.
Denote by $\mathbf{c}_{2}=\nabla I((\widehat{c}_{1},\mathbf{0}))$, which is the value of $\bm{\lambda}$ where the supremum (2) is attained at $\mathbf{x}=(\widehat{c}_{1},\mathbf{0})$. The intuition behind the vector $\mathbf{c}_{2}$ is explained by Figure 1 below. The assumption (A5) is crucial for the non-degeneracy of the jumps. More precisely, (A5) implies that the support of ${\bm{\xi}}$ cannot be contained in a proper subspace of $\mathbb{R}^{d}$ and that ${\bm{\xi}}$ is non-lattice. Both conditions are necessary for the BRW to be able to reach the target ball $B_{x}$. A consequence of (A6) is that $\phi_{\bm{\xi}}$ is well-defined in a neighborhood of $\mathbf{c}_{2}$. We will see in Proposition 21 that (A3) and (A4) together imply (A6) with $\widehat{c}_{1}=c_{1}$.
@@ 1.2 Main results
For a finite set $G$, we use #$\#G$ to denote its cardinality. Denote by $S:=\{\forall n\geqslant 1,\leavevmode\nobreak\ \#V_{n}>0\}$ the survival event that the underlying branching process survives at all times. Since $\rho>1$, we have $p:=\mathbb{P}(S)>0$.
@@
A natural question is whether the $O_{\mathbb{P}}(\log\log x)$ can be replaced by a $O_{\mathbb{P}}(1)$, which we later formulate as Conjecture 1. We discuss a few pieces of evidence that support this conjecture, and further aspects will be discussed in Section 5. First, this is the case for $d=1$ (Proposition 7 below) and for the branching Brownian motion in $\mathbb{R}^{d}$ (Theorem 1 of [54]). A second piece of evidence is the following tightness result on first passage times that is also central to our proof of Theorem 1.
@@
The analog of Theorem 2 for the maximum of BRW traces back to [21] and [41], which assumed uniformly bounded jumps. The case of unbounded jumps was later resolved by [15] and [16] using a different approach. Theorem 2 generalizes these results to the first passage times of BRW in arbitrary dimensions. In this setting, for the jumps, we only require mean zero and assumption (A6). Since (A3) and (A4) together imply (A6), Theorem 2 applies in the spherically symmetric case as well.
@@
In other words, the collection $(\tau_{x}-\widehat{A}(x))/\log x$ converges to 00 in probability as $x\to\infty$.
There are certain instances for non-spherically symmetric jumps where the asymptotic (7) can be improved. Such a class includes product measures and non-degenerate centered elliptical distributions on $\mathbb{R}^{d}$. [^5] We say that a centered random variable ${\bm{\xi}}$ follows a non-degenerate elliptical distribution if ${\bm{\xi}}\overset{\mathrm{law}}{=}TU$ where $U$ is spherically symmetric and $T$ is an invertible linear transformation. These will be addressed in detail in Section 3.2.
@@
[…]
As yet another consequence of Theorems 2 and 3 (while slightly modifying their proofs), the full range of the BRW forms a dense subset of $\mathbb{R}^{d}$, complementing Corollary 2.5 of [43] that the full range of the BBM is dense.
@@
Assume (A1), (A2), (A5), and (A6). Conditioned upon survival, the set

### v2

@@ Abstract
We study the first passage times of discrete-time branching random walks in ${\mathbb{R}}^{d}$ where $d\geqslant 1$. Here, the genealogy of the particles follows a supercritical Galton-Watson process. We provide asymptotics of the first passage times to a ball of radius one at a distance $x$ from the origin, conditioned upon survival. We explicitly provide the linear dominating term and the logarithmic correction term as functions of $x$. The asymptotics are precise up to an order of $o_{\mathbb{P}}(\log x)$ for general jump distributions and up to $O_{\mathbb{P}}(\log\log x)$ for spherically symmetric jumps. A crucial ingredient of both results is the tightness of first passage times. We also discuss an extension of the first passage time analysis to a modified branching random walk model that has been proven to successfully capture shortest path statistics in polymer networks.
first passage time,
keywords:
[class=MSC]
keywords:
[…]
, , and
@@ 1 Introduction
We consider a discrete-time branching random walk (BRW) on $\mathbb{R}^{d}$, where particles in $\mathbb{R}^{d}$ perform independent random walks until they randomly branch into more particles or die. Spatial branching processes, including BRW and its continuous-time sibling branching Brownian motion (BBM), have a long history and a wide range of applications in ecology, population biology, and modeling epidemics ([ermakova2019branching, fisher1937wave, kolmogorov1937etude, konig2020branching, kot2004stochasticity, wang1980convergence]). The study of the extremal behavior of spatial branching processes has gained increasing attention over recent years. The mathematical techniques applied in those studies are related to developments in numerous other fields, e.g., discrete Gaussian free field [bramson2016convergence2, ding2014extreme], random unitary matrices [arguin2017maximum], Riemann zeta function [arguin2017maxima], random multiplicative functions [harper2020moments], spin glasses [bovier2017gaussian], holomorphic multiplicative chaos [gu2024universality], among many others that exhibit a log-correlated structure. We refer to [shi2015branching, zeitouni2016branching] for lecture notes on BRW and related topics.
@@
The study of the extrema for one-dimensional BBM traces back to [bramson1978maximal], which established a precise asymptotic for the maximum as a function of time. For general branching random walks, [addario2009minima] provided a precise asymptotic in terms of the jump distribution of the process. Finer behavior near the frontier is also well understood since the works of [aidekon2013convergence, bramson2016convergence, lalley1987conditional, madaule2017convergence]. Nevertheless, less is known in dimensions greater than one. To name a few examples, there are studies of the maximum norm for BBM ([berestycki2024extremal, kim2023maximum, kim2024shape, mallein2015maximal, stasinski2021derivative]) and very recently of BRW ([bezborodov2023maximal]), as well as the range of spatial branching processes ([kyprianou2002asymptotic, zhang2024large]). These works in higher dimensions typically assumed that the increment distribution is spherically symmetric. A notable exception is [uchiyama1982spatial], which provided estimates on the number of particles in a linearly shifted ball in time for a continuous-time version of the BRW.
@@
Our primary focus is on the first passage times (FPT) for branching random walks in $\mathbb{R}^{d}$. By definition, the FPT is the first time for particles to reach a prescribed region. For BBM, [zhang2024modeling] established a precise asymptotic formula (up to $O_{\mathbb{P}}(1)$) for the FPT; see (62) below. In this paper, we provide FPT asymptotics for BRW that may not have a spherically symmetric jump distribution (while the case of spherically symmetric jumps is of special interest for our applications; see Section 1.4). The main results are explained in Section 1.2 and their proofs are presented in Section 2. Applications to polymer physics will be briefly discussed in Section 1.4 and further detailed in Section 3.1.
We take slightly different routes depending on whether the jump distribution is spherically symmetric or not. In the symmetric case, a key part is establishing asymptotic upper and lower bounds on the number of particles that are near the frontier for a one-dimensional BRW (Proposition 8), where our approach follows the modified second moment method of [bramson2016convergence]. In the non-spherically symmetric case, our argument relies heavily on the theory of random walks in cones [denisov2015random]. In both settings, we prove the tightness around the median of the FPT (Theorem 2) using independent techniques, building upon ideas from [mcdiarmid1995minimal]. A more detailed proof sketch can be found in Section 1.3.
@@ 1.1 Notation and assumptions
Let us start with a formal definition of the branching random walk (BRW). We consider a discrete-time BRW model with offspring distribution $\{p_{i}\}_{i\geqslant 0}$. This means that at each time step, each particle is independently replaced by $i$ particles at the same location with probability $p_{i}$. The particles in each generation have i.i.d. displacements. The case of $i=0$ corresponds to the particle being terminated. Let $\rho=\sum_{i}i\,p_{i}$ be the mean of the offspring distribution. We will always assume the supercritical case $\rho>1$. The $d$-dimensional random walk increment is given by $\bm{\xi}$}, and the first particle starts from the origin $\mathbf{0}\in\mathbb{R}^{d}$. Typically, in this paper, a bold symbol refers to a vector.
Our goal is to study the asymptotics of the FPTs of the BRW. For $x\in\mathbb{R}$, we let $B_{x}$ denote a ball of radius one [^1] Replacing the radius one by any fixed positive constant does not change the analysis; the same proof goes through. centered at $(x,0,\dots,0)$ in $\mathbb{R}^{d}$. We let $V_{n}$ denote the collection (i.e., set) of particle locations at time step $n$, and $\{\bm{\eta}_{v,n}(k)\}_{0\leqslant k\leqslant n}$ denotes the $d$-dimensional random walk that leads to $v\in V_{n}$. We define the FPT $\tau_{x}$ of the BRW to $B_{x}$ as
$\tau_{x}=\min\{n\geqslant 0:\exists\,v\in V_{n},\penalty 10000\ \bm{\eta}_{v,n}(n)\in B_{x}\}.$
In what follows, we will separately analyze the cases depending on whether the law of $\bm{\xi}$} is spherically symmetric or not. [^2] A probability distribution on $\mathbb{R}^{d}$ is spherically symmetric if it is invariant under any orthonormal transformation. Equivalent definitions can be found in Theorem 2.5 of [fang2018symmetric]. In both cases, we impose the following assumptions on the BRW:
@@
For a stochastic process $\{T_{x}\}_{x\geqslant 0}$ and a positive deterministic map $\{t_{x}\}_{x\geqslant 0}$, we write $T_{x}=O_{\mathbb{P}}(t_{x})$ if the collection $\{T_{x}/t_{x}\}_{x\geqslant 0}$ is tight, and $T_{x}=o_{\mathbb{P}}(t_{x})$ if $T_{x}/t_{x}\to 0$ in probability.
@@ 1.1.1 Spherically symmetric jumps
Denote the first coordinate of $\bm{\xi}$} by $\xi^{(1)}$, which is a real-valued random variable. We introduce the large deviation rate function
$I(x):=\sup_{\lambda>0}\Big(\lambda x-\log\phi_{\xi^{(1)}}(\lambda)\Big),$ (1)
where $\phi_{\xi^{(1)}}(\lambda):=\mathbb{E}[e^{\lambda{\xi^{(1)}}}]$ is the moment generating function for $\xi^{(1)}$}. In this case, we make the following assumptions on $\bm{\xi}$}:
@@
    the law of $\bm{\xi}$} is spherically symmetric in $\mathbb{R}^{d}$, and its radial component $R=|{\bm{\xi}}|$ satisfies $\mathbb{P}(R=0)<1$;
@@
    $\log\rho\in(\mathrm{ran}I)^{\circ}$, where $(\mathrm{ran}I)^{\circ}$ is the interior of the range of $I$. In other words, there exists $c_{1}>0$ such that $I(c_{1})=\log\rho$. It can be shown that $c_{1}\in(\mathrm{ran}(\log\phi_{\xi^{(1)}})^{\prime})^{\circ}$.
Note that under assumptions (A2) and (A3), $I$ is $C^{1}$ and convex in its domain, and attains a unique global minimum at x=0x=0.333See e.g., Theorem 26.3 of [rockafellar1970convex]. We will also see from Lemma 24 below that (A3) implies that both $\bm{\xi}$} and $\xi^{(1)}$} are non-lattice if $d\geqslant 2$. [^4] We say the law of $\bm{\xi}$} is non-lattice if there is no $\mathbf{y}\in\mathbb{R}^{d}$ such that the support of ${\bm{\xi}}+\mathbf{y}$ is contained in a discrete subgroup of $\mathbb{R}^{d}$. A consequence of (A4) is that $\rho>1$, and hence the particles form a supercritical Galton-Watson process. Denote by $c_{2}:=I^{\prime}(c_{1})$ the value of $\lambda$ at which the supremum in (1) is taken for $x=c_{1}$. In particular, (A4) implies that $\phi_{\xi^{(1)}}$ is well-defined in a neighborhood of $c_{2}$. The constants $c_{1}$ and $c_{2}$ here depend only on the offspring distribution (through $\rho$) and the increment distribution (through $\phi_{\xi^{(1)}}(\lambda)$) and will be fixed throughout this paper.
@@ 1.1.2 Non-spherically symmetric jumps
Denote by $\phi_{\bm{\xi}}(\bm{\lambda})=\mathbb{E}[e^{\bm{\lambda}\cdot{\bm{\xi}}}]$ the moment generating function for $\bm{\xi}$}. Let
@@
    $\bm{\xi}$} satisfies the strong non-lattice condition $\limsup_{|\mathbf{t}|\to\infty}|\mathbb{E}[e^{i\mathbf{t}\cdot{\bm{\xi}}}]|<1$;
@@
    $\log\rho\in(\mathrm{ran}I(\cdot,\mathbf{0}))^{\circ}$, where $I(\cdot,\mathbf{0})$ refers to the function $I$ with the last d1d-1 variables fixed at zero; let $\widehat{c}_{1}>0$ satisfy $I(\widehat{c}_{1},\mathbf{0})=\log\rho$, then $(\widehat{c}_{1},\mathbf{0})\in(\mathrm{ran}\nabla\log\phi_{\bm{\xi}})^{\circ}$.
Denote by $\mathbf{c}_{2}=\nabla I(\widehat{c}_{1},\mathbf{0})$, which is the value of $\bm{\lambda}$ where the supremum (2) is attained at $\mathbf{x}=(\widehat{c}_{1},\mathbf{0})$. The intuition behind the vector $\mathbf{c}_{2}$ is explained by Figure 1 below. The assumption (A5) is crucial for the non-degeneracy of the jumps. More precisely, (A5) implies that the support of $\bm{\xi}$} cannot be contained in a proper subspace of $\mathbb{R}^{d}$. This is necessary for the BRW to be able to reach the target ball $B_{x}$ and shows that $d$ is the ‘real’ dimension of the increments. A consequence of (A6) is that $\phi_{\bm{\xi}}$ is well-defined in a neighborhood of $\mathbf{c}_{2}$. We will see in Proposition 21 that (A3) and (A4) together imply (A6) with $\widehat{c}_{1}=c_{1}$.
@@ 1.2 Main results
For a finite set $G$, we use #G\#G to denote its cardinality. Denote by $S:=\{\forall n\geqslant 1,\penalty 10000\ \#V_{n}>0\}$ the survival event that the underlying branching process survives at all times. Since $\rho>1$, we have $p:=\mathbb{P}(S)>0$.
@@
A natural question is whether the $O_{\mathbb{P}}(\log\log x)$ can be replaced by a $O_{\mathbb{P}}(1)$, which we later formulate as Conjecture 1.555We remark that the conjecture has been resolved by [blanchet2024tightness] after the first version of this paper was posted; see Remark 4. We discuss a few pieces of evidence that support this conjecture, and further aspects will be discussed in Section 5. First, this is the case for $d=1$ (Proposition 7 below) and for the branching Brownian motion in $\mathbb{R}^{d}$ (Theorem 1 of [zhang2024modeling]). A second piece of evidence is the following tightness result on first passage times that is also central to our proof of Theorem 1.
@@
The analog of Theorem 2 for the maximum of BRW traces back to [dekking1991limit] and [mcdiarmid1995minimal], which assumed uniformly bounded jumps. The case of unbounded jumps was later resolved by [bramson2007tightness] and [bramson2009tightness] using a different approach. Theorem 2 generalizes these results to the first passage times of BRW in arbitrary dimensions. In this setting, for the jumps, we only require mean zero and assumption (A6). Since (A3) and (A4) together imply (A6), Theorem 2 applies in the spherically symmetric case as well.
@@
In other words, the collection $(\tau_{x}-\widehat{A}(x))/\log x$ converges to 0 in probability as $x\to\infty$.
There are certain instances for non-spherically symmetric jumps where the asymptotic (7) can be improved. Such a class includes product measures (meaning that $\bm{\xi}$} has independent coordinates) and non-degenerate centered elliptical distributions on $\mathbb{R}^{d}$. [^6] We say that a centered random variable $\bm{\xi}$} follows a non-degenerate elliptical distribution if ${\bm{\xi}}\overset{\mathrm{law}}{=}TU$ where $U$ is spherically symmetric and $T$ is an invertible linear transformation. These will be addressed in detail in Section 3.2.
@@
$\frac{\tau_{x}}{x}\to\frac{1}{\widehat{c}_{1}}\qquad\text{a.s.}$
As yet another consequence of Theorems 2 and 3 (while slightly modifying their proofs), the full range of the BRW forms a dense subset of $\mathbb{R}^{d}$, complementing Corollary 2.5 of [oz2023] that the full range of the BBM is dense.
@@
Assume (A1), (A2), (A5), and (A6), and that the MGF $\phi_{{\bm{\xi}}}$ is well-defined in a neighborhood of $\mathbf{0}$. Conditioned upon survival, the set